OpenStudy (anonymous):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

OpenStudy (anonymous):

solve by rearranging the equations using algebra and substitution. 2X+Y = 3 2x = 3-y (subtract y from both sides) x= (3-y)/2 (divide both sides by 2) x=2y-1 (second equation) (3-y)/2=2y-1 (substitute from above prob) 3-y=4y-2 (multiply both sides by 2) 3+2=5y (add two and add y to both sides of the equation) 5=5y y=1

OpenStudy (anonymous):

oh, and x=(3-y)/2, so if y=1, then x=(3-1)/2=2/2=1